OpenStudy (anonymous):

The rules for exponents apply to logarithms. In your own words, state the five rules of logarithms and give a unique example to illustrate at least two of the rules.

OpenStudy (inkyvoyd):

log(ab)=log a+log b

OpenStudy (inkyvoyd):

log(a^b)=log a log b

OpenStudy (inkyvoyd):

log(a/b)=log a -log b (I think, may be wrong)

OpenStudy (chaise):

The second law you have posted is incorrect. log(a)^b = b*log(a)

OpenStudy (anonymous):

You can use an example of Inverse properties, product, quotient, or power

OpenStudy (inkyvoyd):

the basic definition of logairtms is also a rule

OpenStudy (inkyvoyd):

and the last rule is the logarithm base changing rule (has to do with division). I just listed them all, so you should be able to do the rest.

OpenStudy (campbell_st):

x^0 = 1 then ln(1) = 0

OpenStudy (campbell_st):

\[10^{\log _{10}x} = x\]

OpenStudy (chaise):

Product: \[\log(a \times b)=\log(a)+\log(b)\] Quotient: \[\log(a-b)=\frac{\log(a)}{\log(b)}\] Power: \[\log(a)^b=b \times \log(a)\] I couldn't be bothered writing out the last. ^_^

OpenStudy (anonymous):

Thanks (: